OpenStudy (anonymous):

Linear equations: -1=5(x-4)-(x-7) What is the solution set? How do I solve this,or is it unsolvable?

OpenStudy (anonymous):

Simplify the right side of your equation first. Then tell me what you get.

OpenStudy (anonymous):

Idl how to simplify it! That's the problem :(

OpenStudy (anonymous):

*idk

OpenStudy (lilg132):

open the brackets up

OpenStudy (amtran_bus):

@lilg132 , why not add one to both sides and set each bracket = to zero?

OpenStudy (anonymous):

How do I combine the two Xs together?

OpenStudy (lilg132):

-1 = 5(x-4)-(x-7) -1 = 5x - 20 - x + 7 now simplify this -1 = 5x - x -20 + 7 -1 = 4x - 13 -1 + 13 = 4x 12 = 4x 4x = 12

OpenStudy (lilg132):

@AmTran_Bus just the way i do it

OpenStudy (anonymous):

So the solution set would be 3?